11-1 Overview of Capital Budgeting-Part 2 Aproject has an internal rate ofreturn of21%, while the company has . wACC of 11%. Should the company go forward with the project? Why? 11-2 NPV Estra Video Determine the net present value of the cash flow shown below ssuming a WACC of 8%. Period 0: -1000 Period 1:400 Period 2: Period 3:3000 Period 4: 100 11-2 Net Present Value The following two projects are mutually exclusive. Based on NPV, what project should the company choose to invest in? Assume each has a WACC of 5%. Time Period Cash Flow for Project I ICash Flow for Project 2 2200 500 400 500 1000 1000 500 500 600 700 800 11-3 IRR YTM A bond is currently selling for $1,300 and pays S65tyear (on an annual basis) for 10 years What is the IRR of the bond? Note that this is the same as asking the bonds YTM) 11-3 Internal Rate of Return-Part 1 Calculate the IRR of the following projeet Time Period Cash Flow -1000 200 300 300 11-4 Multiple Internal Rates of Return
Answer #1

Answer 11-1

A project is acceptable when IRR is greater than required rate of return. So the company should go forward with the project as its IRR(21%) is greater than its WACC(11%).

Answer 11-2

Year Cashflow PVF@8% Cashflow*PVF
0                (1,000) 1                (1,000.00)
1                      400 0.9259                      370.37
2                         -   0.8573                                -  
3                  3,000 0.7938                   2,381.50
4                      100 0.7350                         73.50

NPV = PV of Inflows - PV of Outflows

= (370.37+0+2381.50+73.50) - 1000

= 2825.37 - 1000

= 1825.37

You can use the equation 1/(1+i)^n to find PVF using calculator

Formula to calculate NPV in excel is as follows "=NPV(Rate,cashflows)+initial investment"

Answer 11-2

-Project 1

Year Cashflow PVF@5% Cashflow*PVF
0                (2,200) 1                (2,200.00)
1                      500 0.9524                      476.19
2                      400 0.9070                      362.81
3                      500 0.8638                      431.92
4                  1,000 0.8227                      822.70
5                  2,000 0.7835                   1,567.05

NPV = PV of Inflows - PV of Outflows

= (476.19+362.81+431.92+822.70+1567.05) - 2200

= 3660.68 - 2200

= 1460.68

-Project 2

Year Cashflow PVF@5% Cashflow*PVF
0                (1,000) 1                (1,000.00)
1                      500 0.9524                      476.19
2                      500 0.9070                      453.51
3                      600 0.8638                      518.30
4                      700 0.8227                      575.89
5                      800 0.7835                      626.82

NPV = PV of Inflows - PV of Outflows

= (476.19+453.51+518.30+575.89+626.82) - 1000

= 2650.72 - 1000

= 1650.72

Decision: Project 2 as it has Higher NPV

Answer 11-3

Yield To Maturity(YTM) = (interest per period+ ((Redemption price - Current market price) / life remaining to maturity)) / ((.4*Redemption price)+ (.6*Current market price))

= (65+((1000-1300)/10)) / (.4*1000 + .6*1300)

= (65-30) / 1180

= 35 / 1180

=  0.02966101694

= 2.97%

note: It is general practice to take $1,000 as face value when no details are given.

Answer 11-3

IRR is the rate at which NPV=0. ie: PV of inflows = PV of outflows. It is calculated by trial and error method.

Lets find NPV at say 36%.

Year Cashflow PVF@36% Cashflow*PVF
0                (1,000) 1                (1,000.00)
1                      200 0.7353                      147.06
2                      300 0.5407                      162.20
3                      500 0.3975                      198.77
4                      300 0.2923                         87.69
5                  2,000 0.2149                      429.87

NPV = PV of Inflows - PV of Outflows

= (147.06+162.20+198.77+87.69+429.87) - 1000

= 1025.59-1000

= 25.59

Since NPV is positive, Take a higher rate say 37%

Year Cashflow PVF@37% Cashflow*PVF
0                (1,000) 1                (1,000.00)
1                      200 0.7299                      145.99
2                      300 0.5328                      159.84
3                      500 0.3889                      194.45
4                      300 0.2839                         85.16
5                  2,000 0.2072                      414.41

NPV = PV of Inflows - PV of Outflows

= (145.99+159.84+194.45+85.16+414.41) - 1000

= 999.84-1000

= -.16

Now we got two rates R1 and R2 such that NPV at R1(NPV1) is higher and NPV at R2(NPV2) is lower.

IRR = R1 + ((NPV1 x (R2 - R1)) / (NPV1 - NPV2))

= 36 + ((25.59*(37-36)) / (25.59+.16)

= 36 + (25.59 / 25.75)

= 36 + 0.99378640776

= 36.99%

Actually in this case the IRR comes to 37% in the initial step itself. But I have done further steps only for your understanding.